OpenStudy (anonymous):

what is the answer for the 17th question?

OpenStudy (anonymous):

pls answer

OpenStudy (experimentx):

x^3 + 1/x^3 = (x + 1/x)(x^2 + x*1/x + 1/x^2) = 7 ( 7^2 - 2 + 1) = 7*48

OpenStudy (experimentx):

x^2+1/x^2 = (x+1/x)^2 - 2

OpenStudy (anonymous):

idid nt understand a thing

OpenStudy (experimentx):

a^3+b^3 = (a+b)(a^2-ab+b^2) x^3 + 1/x^3 = (x + 1/x)(x^2 - x*1/x + 1/x^2)

OpenStudy (experimentx):

7*47

OpenStudy (anonymous):

how is 1/x^3 =b^3????

OpenStudy (experimentx):

(1/x)^3 = 1/x^3

OpenStudy (anonymous):

\[(x)^{3}+(1/x)^{3}= (x+1/x) (x ^{2}-x \times1/x+1/x ^{2})\]

OpenStudy (anonymous):

\[= 7( ???????)\]

OpenStudy (experimentx):

x^2+1/x^2 = (x+1/x)^2 - 2

OpenStudy (experimentx):

7^2 - 2

OpenStudy (anonymous):

cn u do it the equation way...itll be more helpful

OpenStudy (experimentx):

\[ x^2+1/x^2 = (x+1/x)^2 - 2 = 49 -2 = 47\]

OpenStudy (anonymous):

thank u

OpenStudy (experimentx):

-x/1/x = -1

OpenStudy (experimentx):

7*46

OpenStudy (anonymous):

wait 1more doubt